# PTR: A Pre-trained Language Model for Trajectory Recovery

### v1

@@ 4.2.2. Implicit Trajectory Prompt
Obviously, the road conditions of a trajectory point are indexed by its longitude, latitude, and timestamp. However, it is challenging for missing points as their geographic coordinates are unknown. To address this, we introduce a road condition diffusion strategy that approximates the missing conditions using neighboring traffic. The implementation details will be elaborated in Section 4.3.2.
@@
4.3. Trajectory Embedder
To efficiently encode trajectory points and enable PLMs to better understand them, we introduce the trajectory embedder.
4.3.1. Trajectory Preprocessing Layer
For different sparse trajectories, their sampling intervals $\mu$ are not constant, such as 4 minutes, 2 minutes, or 1 minute. To convert these into a uniform sampling interval, we introduce a placeholder ‘[m]’ to mark missing trajectory points and generate a preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, where the length of $|\mathcal{T}^{\epsilon}|=\frac{p_{|\mathcal{T}_{s}|}.t-p_{1}.t}{\epsilon}$. Although the location of ‘[m]’ is unknown, its timestamp can still be calculated, as illustrated in Figure 4.
@@ 4.3.2. Trajectory Feature Extractor
To embed trajectory points, we propose the trajectory feature extractor. Given the preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, there are two cases for its trajectory point characteristics:
@@
For case 1, we use Learnable Fourier Features (Tancik et al., 2020; Li et al., 2021) to encode the continuous latitude and longitude of $s$ into an $F$-dimensional vector using the feature mapping function $\Phi(x):\mathbb{R}\to\mathbb{R}^{F}$. Formally:
@@
where $W_{r}\in\mathbb{R}^{F/2}$ and $W_{\Phi}\in\mathbb{R}^{F\times F}$ represent the learnable parameters, and $x\in\{s.lat,s.lon\}$.
Using the feature mapping function $\Phi(\cdot)$, the relative information between points $x$ and $y$ can be captured through multiplication operations. This relative information helps infer crucial details like the distance between trajectory points, which is important for understanding movement. For instance, a larger distance between two points may indicate a higher likelihood of vehicle acceleration.
@@
(5) $\mathbf{h}_{s}^{\text{road}}=\frac{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})*\textbf{M}_{l}}{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})},$
@@
where $W_{1}\in\mathbb{R}^{2F\times F}$ and $b_{1}\in\mathbb{R}^{F}$ is the learnable parameters.
For case 2, where the location of point $s$ is unknown, here we use road conditions to represent its features, as described in Section 4.2.2. Since the road conditions at a specific point are influenced by the surrounding conditions, which propagate along both temporal and spatial dimensions, we propose a road condition diffusion strategy to model the road conditions of the missing location.
First, for the missing point $s$, we identify the observed forward and backward trajectory points, $s_{f}$ and $s_{b}$, respectively. For instance, in Figure 4, suppose we have a missing point $s$ timestamped at 8:00:45, its observed forward point $s_{f}=p_{1}$ and the backward point $s_{b}=p_{2}$. Then, we retrieve the road conditions of $s_{f}$ and $s_{b}$ as follows:
@@
where $\pi_{lat}$, $\pi_{lng}$, and $\pi_{t}$ are index functions mapping latitude, longitude, and time to their respective indices. Next, we calculate the time interval between $s$ and $s_{f}$ and $s_{b}$, denoted as $\triangle t_{f}=s.t-s_{f}.t$ and $\triangle t_{b}=s_{b}.t-s.t$. The road condition of point $s$ is then computed as:
@@
(9) $\begin{split}\mathbf{h}_{s}&=W_{2}[\textbf{m}\text{ }||\text{ FC}(\triangle t_{f}||\triangle t_{b})\text{ }||\text{ }\mathbf{h}_{rc}^{s}]+b_{2},\\ \text{FC}(\cdot)&:\mathbb{R}^{2}\to\mathbb{R}^{F},\end{split}$
where $\textbf{m}\in\mathbb{R}^{F}$ is a learnable vector to represent the location is missing, and FC()($\cdot$)( ⋅ ) is the fully connection layer. $W_{2}\in\mathbb{R}^{3F\times F}$ and $b_{2}\in\mathbb{R}^{F}$ are the learnable parameters. Finally, the overall trajectory representation is $\textbf{H}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times F}$.
@@ 4.4. PLM-based Trajectory Encoder
As shown in Figure 3, we use the stacked Transformer layers as the fundamental architecture of PLMs. To make PLMs fully adaptable to the trajectory recovery task, we employ the Low-Rank Adaptation (LoRA) algorithm (Hu et al., 2021). LoRA introduces additional parameters within the Transformer block, reducing the number of trainable parameters and computational complexity while enhancing the PLM’s performance on new tasks.
@@
After stacking $L$ Transformer layers, we discard the explicit trajectory prompt portion and obtain the output embeddings $\mathbf{Z}^{\prime}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times d}$.
For the $j$-th element of $\textbf{Z}^{\prime}$, we apply the softmax function to calculate the probability of each road segment and use the argmax function to return the predicted road segment $e_{j}$. We employ a MLP with a Sigmoid activation function to determine the moving ratio $r_{j}$.
@@ 4.5.1. Loss Function
We use multi-task learning to simultaneously optimize road segment recovery and moving ratio recovery tasks. For the road segment recovery, we use the cross-entropy as the loss function:
@@
For the moving ratio, we use the mean squared error as the loss function:
@@
4.5.2. Training Strategy
To generalize PTR in recovering sparse trajectories with various sampling intervals, we propose the following training approach.
For the dense trajectory dataset $\mathbb{T}$, which is set as a 15-second sampling interval in our experiments. To create sparse trajectory datasets, we re-sample $\mathbb{T}$ at three different intervals: 1 minute, 2 minutes, and 4 minutes, resulting in three datasets: $\mathbb{T}_{1}$, $\mathbb{T}_{2}$, and $\mathbb{T}_{4}$. These are then combined to form a larger training dataset, $\mathbb{T}_{all}$.
@@ 4.5.2. Training Strategy
By training PTR on $\mathbb{T}_{all}$, we achieve balanced performance across the different sampling intervals. Subsequently, we fine-tune the model on each individual sparse trajectory dataset to further improve trajectory recovery performance.
@@ 5. Experiments
In this section, we present comprehensive experiments to evaluate the effectiveness of PTR on two real-world trajectory datasets with three different sparse trajectory sampling intervals. These experiments are designed to answer the following key questions:
  • RQ1: How does PTR’s performance compare to other baseline methods across different sampling intervals of sparse trajectories?
  • RQ2: How well does PTR perform when trained with limited data?
  • RQ3: What is the contribution of each component in PTR to the overall model performance?
@@ 5.2. Baselines
To evaluate the effectiveness of our model, we compare PTR with ten baseline methods. These include five free space trajectory recovery models: HMM (Newson and Krumm, 2009) + ShortestPath, Linear (Hoteit et al., 2014) + HMM (Newson and Krumm, 2009), MPR (Chen et al., 2011) + HMM (Newson and Krumm, 2009), DHTR (Wang et al., 2019) + HMM (Newson and Krumm, 2009), and AttnMove (Xia et al., 2021) + Rule, as well as five map-matched trajectory recovery models: MTrajRec (Ren et al., 2021), T2vec (Li et al., 2018) + Decoder, T3s (Yang et al., 2021) + Decoder, RNTrajRec (Chen et al., 2023b), and MM-STGED (Wei et al., 2024). The details of each baseline are provided in Appendix B.
@@ 5.3. Evaluation Metrics
We employ five common metrics to assess the effectiveness of our model followed (Ren et al., 2021; Chen et al., 2023b; Wei et al., 2024). For road segment recovery, we utilize Accuracy (Acc), Recall, and Precision (Prec) metrics to evaluate. A higher result indicates a more precise road segment recovery. And employ Mean Absolute Error (MAE) and the Root Mean Square Error (RMSE) in assessing the recovered GPS coordinates. Lower values of these metrics indicate a higher level of accuracy in the recovery results. The details of each metric are provided in Appendix C.
@@ 5.4. Settings
During training, we integrate sparse trajectories with three different sampling intervals to train PTR. After training, we evaluate its performance in recovering trajectories across various sampling intervals to assess its generalizability. To further enhance performance, we fine-tune the trained model for each specific sampling interval, allowing for a more precise recovery.
We employ PyTorch (Paszke et al., 2019) framework to implement PTR, with a learning rate of 1e-4 and a batch size of 64. BERT (Devlin et al., 2018) is selected as the foundation model for PLMs, as its encoder-only structure is well-suited to the reconstruction task due to its effective use of bi-directional contextual information from the trajectory (Ren et al., 2024). For road condition extraction, the area of interest is divided into a $64\times 64$ grid, and the time dimension is partitioned into hourly intervals. The hidden state dimension is set to $F=512$. In Equation 4, we set $\kappa=15$ and $\varphi_{dist}=50$ meters. The model is trained for 50 epochs with early stopping, using a patience of 10 epochs. All experiments are conducted on NVIDIA RTX A4000 GPUs.
Table 2. Scalibility analysis on Chengdu dataset.
Sampling Interval Rate 20% 40% 60% 80% 100%
@@
PTR 76.29 452.8 79.37 411.5 80.37 395.4 81.18 377.4 81.76 366.2
RNTrajRec 75.65 846.2 79.48 784.8 79.61 769.3 81.74 742.5 81.88 702.5
MM-STGED 76.02 857.7 79.86 734.7 82.25 676.4 83.44 663.0 84.26 633.5
PTR 81.62 397.5 84.59 352.0 85.75 328.7 86.73 305.4 87.17 290.8
5.5. Experimental Results (RQ1)
@@
Compared to other baselines, our model achieves superior performance across all metrics, with an average improvement of 16.51% on the Chengdu dataset and 9.35% on the Porto dataset. Notably, it shows substantial gains in RMSE metrics, with an average reduction of 351.8 meters in Chengdu and 119.2 meters in Porto. This indicates PTR generalizes effectively, enabling it to accurately recover trajectories with varying sampling intervals. By fine-tuning at the specific sampling interval, PTR + FT, our model further improves Acc by 1.72% and 0.32%. These outstanding results can be attributed to the strong generalization capability of the PLM in processing sparse trajectories with varying sampling intervals. Additionally, the trajectory embedder plays a pivotal role in connecting trajectory data with the PLM, efficiently capturing spatial-temporal correlations and converting trajectory embeddings into a format that the PLM can effectively process and understand.
We also analyze the convergence rate of PTR across different sampling intervals, as shown in Figure 5 and  6. It demonstrates rapid convergence within just a few epochs on both the Chengdu and Porto datasets. While fine-tuning enhances performance further, the model stabilizes by the 3th epoch, indicating that PTR can deliver satisfactory results within a short training period.
5.6. Scalibility Analysis (RQ2)
To evaluate the effectiveness of PTR in trajectory recovery when dense trajectory data is limited, we conduct scalability experiments on the Chengdu dataset. Specifically, we train the model using subsets of the training set at 20%, 40%, 60%, 80%, and 100%, and evaluate performance on the test set. The results are presented in Table 2.
As the size of the training set increases, we observe performance improvements across all baseline models. Notably, with only 20% of the training data, PTR already surpasses most baselines, trailing only MM-STGED. With 40% of the training data, PTR outperforms all other models, demonstrating its strong generalization ability. This advantage underscores the practicality of PTR in real-world scenarios, where obtaining large-scale dense trajectory data is often challenging.
Table 3. Ablation study on two datasets with sampling intervals at 2 minutes.
@@ 5.6. Scalibility Analysis (RQ2)
Chengdu w/o explicit trajectory prompt 81.04 84.08 88.07 205.8 394.2
w/o implicit trajectory prompt 81.43 84.28 88.14 191.4 371.1
w/o trajectory prompt 80.70 83.61 87.73 225.9 407.4

### v2

@@ 4.2.2. AF-guided Implicit Trajectory Prompt
Obviously, extracting the road conditions of a trajectory point needs its longitude, latitude, and timestamp. However, it is challenging for missing points as their geographic coordinates are unknown. Inspired by the message passing mechanism (Gilmer et al., 2017), we introduce a road condition passing mechanism that approximates the road conditions of missing points by leveraging the surrounding road conditions. The implementation details will be elaborated in Section 4.3.2.
@@
4.3. Interval-aware Trajectory Embedder
To make the model deal with sparse trajectories with different sampling intervals and capture the spatiotemporal correlation, we propose an interval-aware trajectory embedder.
4.3.1. Trajectory Interval Unification
For different sparse trajectories, their sampling intervals $\mu$ are not constant, such as 4 minutes, 2 minutes, or 1 minute, which bring different spatiotemporal correlations between trajectory points. To this end, we convert these into the same sampling interval with the target trajectory by introducing a placeholder ‘[m]’ to mark missing trajectory points and generate a preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, where the length of $\mathcal{T}^{\epsilon}$ is $\frac{p_{|\mathcal{T}_{s}|}.t-p_{1}.t}{\epsilon}+1$. Although the location of ‘[m]’ is unknown, its timestamp can still be calculated, as illustrated in Figure 4.
@@ 4.3.2. Trajectory Feature Extractor
Given the preprocessed sparse trajectory $\mathcal{T}^{\epsilon}$, there are two cases for extract trajectory point features:
@@
For case 1, we use Learnable Fourier Features (LFF) (Tancik et al., 2020; Li et al., 2021) to encode the continuous latitude and longitude of $s$ into an $F$-dimensional vector using the feature mapping function $\Phi(x):\mathbb{R}\to\mathbb{R}^{F}$. Formally:
@@
where $W_{r}\in\mathbb{R}^{F/2}$ and $W_{\Phi}\in\mathbb{R}^{F\times F}$ represent the learnable parameters, and $x\in\{s.lat,s.lon\}$. Using the feature mapping function $\Phi(\cdot)$, the relative information $x-y$ between points $x$ and $y$ can be captured through multiplication operations, which is important for understanding movement, such as distance. We provide detailed insight and proof of LFF in Appendix D.
@@
(5) $\mathbf{h}_{s}^{\text{road}}=\frac{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})\cdot\textbf{M}_{l}}{\sum_{l=1}^{|\mathcal{V}|}f(d_{s,l})},$
@@
where $W_{1}\in\mathbb{R}^{F\times 2F}$ and $b_{1}\in\mathbb{R}^{F}$ is the learnable parameters.
For case 2, where the location of point $s$ is unknown, here we use road conditions to represent its features, as described in Section 4.2.2. Since the road conditions at a specific point are influenced by the surrounding conditions, which propagate along both temporal and spatial dimensions, we propose a road condition passing mechanism to infer the road conditions of the missing location inspired by the message passing mechanism (Gilmer et al., 2017).
Road Condition Passing Mechanism. First, for the missing point $s$, we identify the observed forward and backward trajectory points, $s_{f}$ and $s_{b}$, respectively. For instance, in Figure 4, suppose we have a missing point $s$ timestamped at 8:00:45, its observed forward point $s_{f}=p_{1}$ and the backward point $s_{b}=p_{2}$. Then, we retrieve the road conditions of $s_{f}$ and $s_{b}$ as follows:
@@
where $\pi_{lat}$, $\pi_{lng}$, and $\pi_{t}$ are index functions mapping latitude, longitude, and time to their corresponding indices. Next, we calculate the time interval between $s$ and $s_{f}$ and $s_{b}$, denoted as $\triangle t_{f}=s.t-s_{f}.t$ and $\triangle t_{b}=s_{b}.t-s.t$. The road condition of point $s$ is then computed as:
@@
(9) $\mathbf{h}_{s}=W_{2}[\textbf{m}\text{ }||\text{ FC}(\triangle t_{f}||\triangle t_{b})\text{ }||\text{ }\mathbf{h}_{rc}^{s}]+b_{2},$
where $\textbf{m}\in\mathbb{R}^{F}$ is a learnable vector to represent the location is missing, and $\text{FC}(\cdot):\mathbb{R}^{2}\to\mathbb{R}^{F}$ is the fully connection layer. $W_{2}\in\mathbb{R}^{F\times 3F}$ and $b_{2}\in\mathbb{R}^{F}$ are the learnable parameters. Finally, the overall trajectory representation is $\textbf{H}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times F}$.
@@ 4.4. PLM-based Trajectory Encoder
We use the pre-trained BERT as the fundamental architecture of PLMs, considering that its encoder-only structure is well-suited to the reconstruction task due to its effective use of bi-directional contextual information from the trajectory (Sun et al., 2023). To make PLM fully adaptable to the trajectory recovery task, we employ the Low-Rank Adaptation (LoRA) algorithm (Hu et al., 2021). LoRA introduces additional parameters within the Transformer block, reducing the number of trainable parameters and computational complexity while enhancing the PLM’s performance on new tasks.
@@
After stacking $L$ Transformer layers, we discard the IF-guided explicit trajectory prompt portion and obtain the output embeddings $\mathbf{Z}^{\prime}\in\mathbb{R}^{|\mathcal{T}^{\epsilon}|\times d}$. For the $j$-th element of $\textbf{Z}^{\prime}$, we apply the softmax function to calculate the probability of each road segment and use the argmax function to return the predicted road segment $e_{j}$. We employ a MLP with a Sigmoid activation function to determine the moving ratio $r_{j}$.
@@ 4.5.1. Loss Function
We use multi-task learning to optimize both road segment recovery and moving ratio recovery simultaneously. For the road segment, we use the cross-entropy loss function:
@@
For the moving ratio, we use the mean squared error loss function:
@@
4.5.2. Joint Training Strategy
To generalize PLMTrajRec in recovering sparse trajectories with various sampling intervals, we propose the following training approach.
For the dense trajectory dataset $\mathbb{T}$, which is set as a 15-second sampling interval in our experiments. To create sparse trajectory datasets, we re-sample each trajectory $\mathcal{T}\in\mathbb{T}$ at three different intervals: 1 minute, 2 minutes, and 4 minutes, resulting in three datasets: $\mathbb{T}_{1}$, $\mathbb{T}_{2}$, and $\mathbb{T}_{4}$. These are then combined to form a larger training dataset $\mathbb{T}_{all}$. Formally,
@@ 4.5.2. Joint Training Strategy
By training PLMTrajRec on $\mathbb{T}_{all}$, we achieve balanced performance across the different sampling intervals. Subsequently, we fine-tune the model on each individual sparse trajectory dataset to further improve trajectory recovery performance.
@@ 5. Experiments
In this section, we present comprehensive experiments to evaluate the effectiveness, scalability, and generalization capabilities of PLMTrajRec on two real-world trajectory datasets with three different sparse trajectory sampling intervals.
@@ 5.1. Datasets
Table 1. Dataset Description.
Types Chengdu Porto
Sampling interval 15s 15s
# Trajectory 118,354 322,079
# Road segment 2504 2224
Latitude range 30.655 $\sim$ 30.727 41.142 $\sim$ 41.174
Longitude range 104.043 $\sim$ 104.129 -8.652 $\sim$ -8.578
Length / Width 8.22 km / 8.00 km 6.19 km / 3.56 km
@@ 5.2. Baselines
To evaluate the effectiveness of our model, we compare PLMTrajRec with 12 baseline methods. These include five free space trajectory recovery models: HMM (Newson and Krumm, 2009) + ShortestPath, Linear (Hoteit et al., 2014) + HMM (Newson and Krumm, 2009), MPR (Chen et al., 2011) + HMM (Newson and Krumm, 2009), DHTR (Wang et al., 2019) + HMM (Newson and Krumm, 2009), and AttnMove (Xia et al., 2021) + Rule, as well as seven map-matched trajectory recovery models: MTrajRec (Ren et al., 2021), T2vec (Li et al., 2018) + Decoder, T3s (Yang et al., 2021) + Decoder, TERI (Chen et al., 2023a) + Decoder, TrajBERT (Si et al., 2023) + Decoder, RNTrajRec (Chen et al., 2023b), and MM-STGED (Wei et al., 2024). Due to space constraints, the details of each baseline are provided in Appendix A.
@@ 5.3. Evaluation Metrics
We employ five common metrics to assess the effectiveness of our model followed (Ren et al., 2021; Chen et al., 2023b; Wei et al., 2024). For road segment recovery, we utilize Accuracy (Acc), Recall, and Precision (Prec) metrics to evaluate. And employ Mean Absolute Error (MAE) and the Root Mean Square Error (RMSE) in assessing the recovered GPS coordinates. The details of each metric are provided in Appendix B.
@@ 5.4. Settings
During training, we integrate sparse trajectories with three different sampling intervals to train PLMTrajRec. After training, we evaluate its performance in recovering trajectories across various sampling intervals to assess its generalizability. To further enhance performance, we fine-tune the trained model for each specific sampling interval, allowing for a more precise recovery.
We employ PyTorch (Paszke et al., 2019) framework to implement PLMTrajRec, with a learning rate of 1e-4 and a batch size of 64. BERT-small  [^4] https://github.com/google-research/bert is selected as the foundation model for PLM with 4 transformer layers and the number of hidden state is 512. For road condition extraction, the area of interest is divided into a $64\times 64$ grid, and the time dimension is partitioned into hourly intervals. The hidden state dimension is set to $F=512$. In Equation 4, we set $\kappa=15$ and $\varphi_{dist}=50$ meters. The model is trained for 50 epochs with early stopping, using a patience of 10 epochs. All experiments are conducted on NVIDIA RTX A4000 GPUs.
Table 3. Scalability analysis. The performance comparison on the Chengdu dataset when trained with different data ratios.
Sampling Interval Data Ratio 20% 40% 60% 80% 100%
@@
PLMTrajRec 76.29 452.8 79.37 411.5 80.37 395.4 81.18 377.4 81.76 366.2
5.5. Experimental Results
@@
Compared to other baselines, our model achieves superior performance across all metrics, with an average improvement of 16.51% on the Chengdu dataset and 9.35% on the Porto dataset. Notably, it shows substantial gains in RMSE metrics, with an average reduction of 351.8 meters in Chengdu and 119.2 meters in Porto. This indicates PLMTrajRec generalizes effectively, enabling it to accurately recover trajectories with varying sampling intervals. By fine-tuning at the specific sampling interval, PLMTrajRec + FT, our model further improves Acc by 1.72% and 0.32%. These outstanding results can be attributed to the strong generalization capability of the PLM in processing sparse trajectories with varying sampling intervals. Additionally, the interval-aware trajectory embedder plays a pivotal role to capture spatial-temporal correlations and converting trajectory embeddings into a format that the PLM can effectively process and understand.
5.6. Scalability Analysis
To evaluate the effectiveness of PLMTrajRec in trajectory recovery when dense trajectory data is limited, we conduct scalability experiments on the Chengdu dataset. Specifically, we train the model using subsets of the training set at 20%, 40%, 60%, 80%, and 100%, and evaluate performance on the test set. The results are presented in Table 3 and 7 of Appendix F.2. As the size of the training set increases, we observe performance improvements across all baseline models. Notably, with only 20% of the training data, PLMTrajRec already surpasses most baselines, trailing only MM-STGED. With 40% of the training data, PLMTrajRec outperforms all other models, demonstrating its strong scalability. This advantage underscores the practicality of PLMTrajRec in real-world scenarios, where obtaining large-scale dense trajectory data is often challenging.
Table 4. Zero-shot study on Chengdu dataset with sampling intervals at 2 minutes.
@@ 5.6. Scalability Analysis
MTrajRec 56.39 ($\downarrow$ 18.13%) 492.7 ($\downarrow$ 48.37%) 1046.2 ($\downarrow$ 15.34%)
T3s + Decoder 63.81 ($\downarrow$ 10.81%) 353.0 ($\downarrow$ 31.39%) 980.6 ($\downarrow$ 9.68%)
T2vec + Decoder 63.32 ($\downarrow$ 12.37%) 376.6 ($\downarrow$ 38.50%) 928.1 ($\downarrow$ 15.57%)
RNTrajRec 64.93 ($\downarrow$ 10.87%) 339.6 ($\downarrow$ 35.70%) 912.8 ($\downarrow$ 17.07%)
MM-STGED 68.27 ($\downarrow$ 9.90%) 291.8 ($\downarrow$ 32.42%) 862.6 ($\downarrow$ 19.31%)
PLMTrajRec 79.62 ($\downarrow$ 2.14%) 195.3 ($\downarrow$ 4.05%) 379.6 ($\downarrow$ 3.53 %)
5.7. Zero-shot Study on Sampling Interval
We further analyze the generalizability of our model when the target sampling interval is absent during training. Specifically, PLMTrajRec is trained with a mixture of sampling intervals of 1 minute and 4 minutes and tested on a 2-minute sampling interval. The experimental results on the Chengdu dataset are shown in Table 4, where $\downarrow$ indicates the percentage decline compared to training with a 2-minute sampling interval. We observe that the accuracy of all models decreased when the target sampling interval was not included in the training data. However, our model shows a notably smaller performance drop. This robustness is attributed to the trajectory prompts and joint training strategy used by PLMTrajRec, which effectively identify the sampling interval of the target trajectories, demonstrating strong generalizability. In contrast, baseline models rely solely on trajectory point information to capture correlations, which struggles with generalization due to the inconsistency of temporal dynamics between the training and testing.
5.8. Ablation Study
To evaluate the effectiveness of the proposed components, we create six variants of PLMTrajRec. w/o IF-guided explicit trajectory prompt, w/o AF-guided implicit trajectory prompt, w/o dual trajectory prompt, w/o road network, w/o reference tokens, and PLMTrajRec - Randomly initialized BERT. The detailed description of each variant is shown in Appendix E.
Table 5. Ablation study on Chengdu dataset with sampling intervals at 2 minutes.
[…]
w/o IF-guided explicit trajectory prompt 81.04 84.08 88.07 205.8 394.2
w/o AF-guided implicit trajectory prompt 81.43 84.28 88.14 191.4 371.1
w/o dual trajectory prompts 80.70 83.61 87.73 225.9 407.4